How would I be able to find the surface area of a hockey rink that is 100 feet by 200 feet, with quarter cirlces in the corners that have a radius of 20 feet?
Thank you very, very much!!! :-)

Answer by checkley71(8403)

The rink is 100*200=20,000ft^2
The 4 quarter circles in the corners equals one circle with a radius=20.
Thus we have to subtract one circle with radius=20 ft
area=pi*r^2
area=3.14*20^2
area=3.14*400
area=1,256
20,000-1,256=18,744 ft^2 is the remaining area of the rink.